OpenStudy (anonymous):

writing an equation of a line that is perpendicular to y=2x-5 and passes through (1,7)

OpenStudy (anonymous):

if i line is perpendicular, that means that the slope is the negative reciprocal. since the slope is 2, the negative reciprocal of 2 is -1/2 plug 1 in for the x coordinate and 7 in for the y 7 = -1/2(1) + b 7 = -1/2 + b b = 7.5 y = -1/2x + 7.5

OpenStudy (anonymous):

i think :)

OpenStudy (anonymous):

thanks :)